# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 5. A derived Milnor-Moore theorem
$\mathrm{colim}:\mathrm{Fun}({\mathbb{N}},{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},$
$\oplus:\prod_{{\mathbb{N}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
are symmetric monoidal with respect to Day-convolution, where ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ carries the cartesian structure and $\mathrm{Fun}({\mathbb{N}},{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}),\prod_{{\mathbb{N}}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ carry the induced Day-convolution symmetric monoidal structures. Hence it is enough to see that the functor
[…]
@@
$\mathrm{Fun}({\mathbb{N}},{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})$
the Day-convolution. By [13, Theorem 1.13.] the functor $\gamma$ forming the associated graded, and so its restriction $\gamma^{\prime}$ are symmetric monoidal with respect to Day-convolution. Because $\oplus\circ\gamma^{\prime}$ is conservative, it is enough to see that the functor $\rho$ preserves finite products. Since the forgetful functor $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathcal{C}}$ preserves finite products, it suffices to see that $\mathrm{Sym}:{\mathcal{C}}\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ preserves finite products. By [23, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ is cartesian.
@@
Next we construct the enveloping Hopf algebra (Definition 5.12). For that we use the notions of monoid object and ${\mathbb{E}}_{\infty}$-monoid object [23, Definition 2.4.2.1., Definition 4.1.2.5.] as well as the notion of group object in any $\infty$-category ${\mathcal{D}}$ that admits finite products. A group object in ${\mathcal{D}}$ is a monoid object having the property that the shearing morphism $(\mu,\mathrm{pr}_{1}):{\mathrm{X}}\times{\mathrm{X}}\to{\mathrm{X}}\times{\mathrm{X}}$ is an equivalence, where $\mu$ is the multiplication of ${\mathrm{X}}$ and $\mathrm{pr}_{1}$ is projection to the first factor.
Notation 5.5.
@@
  • Let $\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\subset\mathrm{Fun}({\mathrm{{\mathcal{F}}in}}_{*},{\mathcal{D}})$ be the full subcategory spanned by the ${\mathbb{E}}_{\infty}$-monoid objects in ${\mathcal{D}}$.
@@
  • Let ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\subset\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ be the full subcategory spanned by the ${\mathbb{E}}_{\infty}$-monoid objects in ${\mathcal{D}}$ whose underlying monoid object in ${\mathcal{D}}$ is a group object.
@@
Let ${\mathcal{D}}$ be an additive $\infty$-category. By definition of additivity ${\mathrm{Grp}}({\mathcal{D}})=\mathrm{Mon}({\mathcal{D}})$ and so ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})=\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}).$ By [23, Proposition 2.4.2.5., Proposition 2.4.3.9.] the forgetful functors ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})=\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\to{\mathcal{D}},{\mathrm{Grp}}({\mathcal{D}})=\mathrm{Mon}({\mathcal{D}})\to{\mathcal{D}}$ are equivalences.
@@
The adjunction ${\mathrm{Bar}}:{\mathrm{Grp}}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ is an equivalence.
@@
As ${\mathcal{C}}$ is additive, by Remark 5.6 the forgetful functors $\mathrm{Mon}({\mathcal{C}})\to{\mathcal{C}},\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ are equivalences. We first assume that ${\mathcal{C}}$ admits geometric realizations and totalizations. In this case by [23, Corollary 3.4.1.7.] there is a canonical equivalence $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathrm{Mod}}_{\mathrm{A}}({\mathcal{C}}))\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathrm{A}}/}\simeq{\mathcal{C}}_{{\mathrm{A}}/}.$ So for any ${\mathrm{A}}\in{\mathcal{C}}\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ the relative tensor product ${\mathrm{Bar}}({\mathrm{A}})=0\otimes_{\mathrm{A}}0$ is the coproduct of ${\mathrm{A}}\to 0$ with itself in the $\infty$-category $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathrm{Mod}}_{\mathrm{A}}({\mathcal{C}}))\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathrm{A}}/}\simeq{\mathcal{C}}_{{\mathrm{A}}/}$ and so identifies with $\Sigma({\mathrm{A}}).$ So the adjunction ${\mathrm{Bar}}:\mathrm{Mon}({\mathcal{C}})\simeq{\mathcal{C}}\rightleftarrows{\mathcal{C}}_{\ast}\simeq{\mathcal{C}}:{\mathrm{Cobar}}$ identifies with the equivalence $\Sigma:{\mathcal{C}}\simeq{\mathcal{C}}:\Omega.$
For general ${\mathcal{C}}$ we embed ${\mathcal{C}}$ exact into the stable presentable $\infty$-category ${\mathrm{Ind}}({\mathcal{C}})$ and find that for every ${\mathrm{A}}\in\mathrm{Mon}({\mathcal{C}})\simeq{\mathcal{C}}$ the simplicial object ${\mathrm{Bar}}(0,{\mathrm{A}},0):\Delta^{\mathrm{op}}\to{\mathcal{C}},[{\mathrm{n}}]\mapsto{\mathrm{A}}^{\oplus{\mathrm{n}}}$ admits a colimit that is given by $\Sigma({\mathrm{A}})$ and for every ${\mathrm{B}}\in{\mathcal{C}}_{*}\simeq{\mathcal{C}}$ the cosimplicial object ${\mathrm{Cobar}}(0,{\mathrm{B}},0):\Delta\to{\mathcal{C}},[{\mathrm{n}}]\mapsto{\mathrm{B}}^{\oplus{\mathrm{n}}}$ admits a limit that is given by $\Omega({\mathrm{B}})$. Because ${\mathcal{D}}$ admits geometric realizations and totalizations, there is an adjunction ${\mathrm{Bar}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}.$ Since $\phi:{\mathcal{D}}\to{\mathcal{C}}$ preserves geometric realizations, finite products and totalizations, $\phi$ sends the unit and counit of the adjunction ${\mathrm{Bar}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ to the unit and counit, respectively, of the adjunction $\Sigma:{\mathcal{C}}\rightleftarrows{\mathcal{C}}:\Omega.$ As $\phi$ is conservative, both adjoints of the adjunction ${\mathrm{Bar}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ are equivalences.
@@
${\mathrm{Bar}}:{\mathrm{Grp}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}))\rightleftarrows\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}):{\mathrm{Cobar}}$
@@
Notation 5.9.
@@
$\mathrm{Bialg}({\mathcal{C}})=\mathrm{Mon}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})),$
${\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
Construction 5.10.
@@
${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}}):{\mathrm{triv}}*^{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}(-).$
@@
${\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},\ {\mathrm{X}}\mapsto{\mathrm{X}}\oplus{\mathbb{1}}.$
@@
${\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},\ {\mathrm{X}}\mapsto{\mathrm{X}}\oplus{\mathbb{1}}$
@@
(12) ${\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))=\mathrm{Mon}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\rightleftarrows\mathrm{Bialg}({\mathcal{C}})\simeq\mathrm{Mon}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}),$
where the left adjoint lands in ${\mathrm{Hopf}}({\mathcal{C}})\simeq{\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}).$
@@
(13) ${\mathrm{Cobar}}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))=\mathrm{Mon}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})):{\mathrm{Bar}}$
covering $\Sigma:{\mathcal{C}}\to{\mathcal{C}}$. Composing adjunctions (13) and (12) gives an adjunction
${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\rightleftarrows\mathrm{Bialg}({\mathcal{C}}):{\mathrm{Prim}}^{\prime},$
where the left adjoint lands in ${\mathrm{Hopf}}({\mathcal{C}}).$
@@
$\mathrm{Bialg}({\mathcal{C}})\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}*^{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}(-)}{\mathcal{C}}\xrightarrow{\Sigma}{\mathcal{C}}.$
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\rightleftarrows\mathrm{Bialg}({\mathcal{C}}):{\mathrm{Prim}}$
@@
${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows\mathrm{Bialg}({\mathcal{C}}):{\mathrm{Prim}}^{\prime}.$
Note that the left adjoint ${\mathcal{U}}$ lands in ${\mathrm{Hopf}}({\mathcal{C}})\subset\mathrm{Bialg}({\mathcal{C}}).$
@@
Since limits in group objects are formed underlying, Proposition 5.4 implies that the functor ${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\to\mathrm{Bialg}({\mathcal{C}})$ preserves finite products and so turns the direct sum to the tensor product.
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\rightleftarrows\mathrm{Bialg}({\mathcal{C}}):{\mathrm{Prim}}$
@@
[…]
@@
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}}{\mathrm{coAlg}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\mathrm{free}}\mathrm{Bialg}({\mathcal{C}})=\mathrm{Mon}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$

### v2

@@ 4. A derived Milnor-Moore theorem
$\mathrm{colim}:\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},$
$\oplus:\prod_{{\mathbb{N}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
are symmetric monoidal with respect to Day-convolution, where $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ carries the cartesian structure and $\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}),\prod_{{\mathbb{N}}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ carry the induced Day-convolution symmetric monoidal structures. Hence it is enough to see that the functor
$\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\xrightarrow{\theta}\mathrm{Fun}({\mathbb{N}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))_{*}}\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})$
@@
$\mathrm{Fun}({\mathbb{N}},\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/})$
the Day-convolution. By [20, Theorem 1.13.] the functor $\gamma$ forming the associated graded, and so its restriction $\gamma^{\prime}$ are symmetric monoidal with respect to Day-convolution. Because $\oplus\circ\gamma^{\prime}$ is conservative, it is enough to see that the functor $\rho$ preserves finite products. Since the forgetful functor $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\to{\mathcal{C}}$ preserves finite products, it suffices to see that ${\mathrm{Sym}}:{\mathcal{C}}\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ preserves finite products. By [39, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$ is cartesian.
@@
[…]
Notation 4.5.
@@
[…]
@@
[…]
@@
Let $\mathcal{D}$} be an additive $\infty$-category. By definition of additivity ${\mathrm{Grp}}({\mathcal{D}})=\mathrm{Mon}({\mathcal{D}})$ and so ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})=\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}).$ By [39, Proposition 2.4.2.5., Proposition 2.4.3.9.] the forgetful functors ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})=\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\to{\mathcal{D}},{\mathrm{Grp}}({\mathcal{D}})=\mathrm{Mon}({\mathcal{D}})\to{\mathcal{D}}$ are equivalences.
@@
The adjunction ${\mathrm{Barc}}:{\mathrm{Grp}}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ is an equivalence.
@@
As $\mathcal{C}$} is additive, by Remark 4.6 the forgetful functors $\mathrm{Mon}({\mathcal{C}})\to{\mathcal{C}},\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ are equivalences. We first assume that $\mathcal{C}$} admits geometric realizations and totalizations. In this case by [39, Corollary 3.4.1.7.] there is a canonical equivalence $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathrm{Mod}}_{\mathrm{A}}({\mathcal{C}}))\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathrm{A}}/}\simeq{\mathcal{C}}_{{\mathrm{A}}/}.$ So for any ${\mathrm{A}}\in{\mathcal{C}}\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ the relative tensor product ${\mathrm{Barc}}({\mathrm{A}})=0\otimes_{\mathrm{A}}0$ is the coproduct of ${\mathrm{A}}\to 0$ with itself in the $\infty$-category $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathrm{Mod}}_{\mathrm{A}}({\mathcal{C}}))\simeq\mathrm{Mon}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathrm{A}}/}\simeq{\mathcal{C}}_{{\mathrm{A}}/}$ and so identifies with $\Sigma({\mathrm{A}}).$ So the adjunction ${\mathrm{Barc}}:\mathrm{Mon}({\mathcal{C}})\simeq{\mathcal{C}}\rightleftarrows{\mathcal{C}}_{\ast}\simeq{\mathcal{C}}:{\mathrm{Cobar}}$ identifies with the equivalence $\Sigma:{\mathcal{C}}\simeq{\mathcal{C}}:\Omega.$
For general $\mathcal{C}$} we embed $\mathcal{C}$} exact into the stable presentable $\infty$-category ${\mathrm{Ind}}({\mathcal{C}})$ and find that for every ${\mathrm{A}}\in\mathrm{Mon}({\mathcal{C}})\simeq{\mathcal{C}}$ the simplicial object ${\mathrm{Barc}}(0,{\mathrm{A}},0):\Delta^{\mathrm{op}}\to{\mathcal{C}},[{n}]\mapsto{\mathrm{A}}^{\oplus{n}}$ admits a colimit that is given by $\Sigma({\mathrm{A}})$ and for every ${\mathrm{B}}\in{\mathcal{C}}_{*}\simeq{\mathcal{C}}$ the cosimplicial object ${\mathrm{Cobar}}(0,{\mathrm{B}},0):\Delta\to{\mathcal{C}},[{n}]\mapsto{\mathrm{B}}^{\oplus{n}}$ admits a limit that is given by $\Omega({\mathrm{B}})$. Because $\mathcal{D}$} admits geometric realizations and totalizations, there is an adjunction ${\mathrm{Barc}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}.$ Since $\phi:{\mathcal{D}}\to{\mathcal{C}}$ preserves geometric realizations, finite products and totalizations, $\phi$ sends the unit and counit of the adjunction ${\mathrm{Barc}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ to the unit and counit, respectively, of the adjunction $\Sigma:{\mathcal{C}}\rightleftarrows{\mathcal{C}}:\Omega.$ As $\phi$ is conservative, both adjoints of the adjunction ${\mathrm{Barc}}:\mathrm{Mon}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ are equivalences.
@@
${\mathrm{Barc}}:{\mathrm{Grp}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}))\rightleftarrows\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}):{\mathrm{Cobar}}$
@@
Corollary 4.9.
$\Omega\circ{\mathrm{Barc}}:{\mathrm{Grp}}(\mathrm{Alg}_{{\mathcal{O}}(1)}({\mathcal{C}}))\rightleftarrows\mathrm{Alg}_{{\mathcal{O}}(1)}({\mathcal{C}})\simeq\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}):{\mathrm{Cobar}}\circ\Sigma$
is an equivalence over $\mathcal{C}$} via the forgetful functors.
Notation 4.10.
@@
${\mathrm{Bialg}}({\mathcal{C}})=\mathrm{Mon}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})),$
${\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$
Remark 4.6 gives the following:
Corollary 4.11.
Let $\mathcal{C}$} be an additive presentably symmetric monoidal $\infty$-category. Then
${\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Bialg}}({\mathcal{C}}).$
Construction 4.12.
@@
${\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-):\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}}):{\mathrm{triv}}*^{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}(-).$
@@
$\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},\ {\mathrm{X}}\mapsto{\mathrm{X}}\oplus{\mathbb{1}}.$
@@
$\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/},\ {\mathrm{X}}\mapsto{\mathrm{X}}\oplus{\mathbb{1}}$
@@
(4.1) ${\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))=\mathrm{Mon}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\rightleftarrows{\mathrm{Bialg}}({\mathcal{C}})\simeq\mathrm{Mon}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}),$
where the left adjoint lands in ${\mathrm{Hopf}}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}).$
@@
(4.2) ${\mathrm{Cobar}}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))=\mathrm{Mon}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})):{\mathrm{Barc}}$
covering $\Sigma:{\mathcal{C}}\to{\mathcal{C}}$. Composing adjunctions (4.2) and (4.1) gives an adjunction
${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\rightleftarrows{\mathrm{Bialg}}({\mathcal{C}}):{\mathrm{Prim}}^{\prime}.$
@@
${\mathrm{Bialg}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}*^{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}(-)}{\mathcal{C}}\xrightarrow{\Sigma}{\mathcal{C}}.$
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Bialg}}({\mathcal{C}}):{\mathrm{Prim}}$
@@
${\mathcal{U}}^{\prime}:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}^{\prime}.$
@@
Since limits in group objects are formed underlying, Proposition 4.4 implies that the functor ${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\to{\mathrm{Hopf}}({\mathcal{C}})$ preserves finite products and so turns the direct sum to the tensor product.
@@
${\mathcal{U}}:\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}$
@@
${\mathrm{Hopf}}({\mathcal{C}})\xrightarrow{\text{forget}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}*^{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}(-)}{\mathcal{C}}.$
@@
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}}\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nun}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\mathrm{free}}{\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Bialg}}({\mathcal{C}})=\mathrm{Mon}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})).$